OpenStudy (anonymous):

Use mathematical induction to prove the statement is true for all positive integers n, or show why it is false. 12 + 42 + 72 + ... + (3n - 2)2 = (n(6n^2-3n-1))/2

OpenStudy (anonymous):

@midhun.madhu1987 If I do this one, can you just check my work over?

OpenStudy (midhun.madhu1987):

yes.. sure.. i can check.. :)

OpenStudy (anonymous):

Awesome, so first I have to square the whole left side, right?

OpenStudy (midhun.madhu1987):

prove for p(1)

OpenStudy (midhun.madhu1987):

The statement is valid only if P(1) is TRUE

OpenStudy (anonymous):

This is what I got: 9n^2-12n+4=(n(6n^2-3n-1))/2 (9(1))^2-12(1)+4=(1((6(1))^2-3(1)-1))/2 9^2-12+4=(1(6^2)-3-1)/2 81-12+4=(36-3-1)/2 73=32/2 73=16 So I'm thinking it's false

OpenStudy (midhun.madhu1987):

 so p(1) is true